A block of glass with refractive index,w.5 is surrounded by air. A rat of light enters the block at its left-hand face with incident angle X and reemerges into the air from the right-hand face directed parallel to the block's base. Find X


1
Expert's answer
2021-10-21T16:33:40-0400

"\frac{\sin y}{\sin x}=n,"

"y=90\u00b0," "\sin y=90°,"

"\frac 1{\sin x}=n,\implies"

"x=\arcsin\frac 1n=41.8°."